if 4 is added to twice a number and the sum is multiplied by 2, the result is the same as if the number is multiplied by 3 and 4 is added to the product. what is the number?
Answer by oberobic(2304)

A number = x.
4 is added to twice a number = 2x + 4
4 is added to three times a number = 3x + 4
2x + 4 = y
3x + 4 = y
.
By the transitive property...
.
2x + 4 = 3x + 4
.
0 = x
.
x = 0
.
So the answer is x = 0.
.
Any number times 0 = 0, so that's logically the only possible answer.
.
Done